Consider a firm in a perfectly competitive market which chooses an input π‘₯ to maximize its profit πœ‹ = 100𝑓(π‘₯) βˆ’ 𝑀π‘₯.Β 

i) Specify the first order condition for the profit maximization problem. (4 marks) ii) Suppose there is a differentiable function π‘₯(𝑀)that satisfies the implicit function in part (i) above. Using the chain rule, find π‘₯β€²(𝑀). (4 marks) iii) Suggest conditions on the function 𝑓(π‘₯) so that π‘₯ β€²(𝑀) < 0



1
Expert's answer
2021-01-14T09:24:52-0500

i)


"\frac{\delta\pi}{\delta x}=100 f'(x)-w"

ii)


"\frac {\delta \pi}{\delta x}=100f'(x)-xw'(x)-w"

iii)


"w'(x)<0"

"w'(x)=\frac{100f'(x)-w(x)}{x}"